OpenStudy (anonymous):

A person with a mass of 80 kg applies a vertical force to the ground of magnitude 2000 N, and has a braking force of 1000 N. What is the resultant acceleration of the person (magnitude only)? (a) 19.7 ms-2 (b) 30.0 ms-2 (c) 28.0 ms-2 (d) 37.0 ms-2 (e) 48.9 ms-2

OpenStudy (anonymous):

is this braking force horizontal? do you know?

OpenStudy (anonymous):

yeah, i'd say it's a horizontal force.

OpenStudy (anonymous):

ok so Lets look at all the vertical forces acting on him Force of gravity F = mg F = 80 x 10 = 800N Vertical force that he exerts 2000 N Sum of all the vert forces = 800 + 2000 = 2800N Now like you said assuming that the braking force is acting horizontally  Using Pythagoras F net = 2973 N F net = m x a to find the resultant acceleration, rearrange for A a = Fnet / m a = 2973 / 80 = 37 ms ^ - 2